Company XYZ prepares this for jam, which are labeled to have a mass of



largebatch containing 100 and is found to have a mass of 500 grams.



Assuming the masses are normally distributed, test whether the mass Of this



550 grams with a standard deviation of 20 grams. One tin is selected from a



500 grams. During the preparation the tins are set to have a mean mass of



tin is significantly different from 550 gram at 5% level of significance

1
Expert's answer
2022-01-17T14:05:18-0500

The hypotheses to be tested are,

"H_0:\mu=550\\
vs\\
H_1:\mu\not=550"

"x=500,\sigma=20,\alpha=0.05"

The test statistic is given as,

"Z={x-\mu\over\sigma}\\
Z={500-550\over20}={-50\over20}=-2.5"

"Z" is compared with the standard normal table value at "\alpha=0.05"

The table value is given as,

"Z_{\alpha\over2}=Z_{0.05\over2}=Z_{0.025}=-1.96"

The null hypothesis is rejected if "Z\lt Z_{0.025}"

Since "Z=-2.5\lt Z_{0.025}=-1.96", we reject the null hypothesis and conclude that there is sufficient evidence to show that the mass of the selected tin is significantly

different from 550 grams at 5% level of significance.